1.A cyclist negotiates a curve of radius 100m with a velocity 10 m/s. How much should he

lean from the vertical in doing so?

2.. A disc revolves in a horizontal plane at a steady rate of 3 revs/s. A coin will remain on the

disc if it is kept a distance of 2 cm from the axis of rotation. What is the coefficient of

friction between the coin and the disc?


1
Expert's answer
2021-05-30T13:30:07-0400

1) Let's apply Newton's Second Law of Motion in projections on axis "x"- and "y":


"Ncos\theta=mg,""Nsin\theta=\dfrac{mv^2}{r}."

Dividing the second equation by the first one, we get:


"tan\theta=\dfrac{v^2}{rg},""\theta=tan^{-1}(\dfrac{v^2}{rg}),""\theta=tan^{-1}(\dfrac{(10\ \dfrac{m}{s})^2}{100\ m\cdot9.8\ \dfrac{m}{s^2}})=5.8^{\circ}."

2)

"F_c=F_{fr},""\dfrac{mv^2}{r}=\mu mg,""\omega^2r=\mu g,""\mu=\dfrac{\omega^2r}{g},""\mu=\dfrac{(3\ \dfrac{rev}{s}\cdot\dfrac{2\pi\ rad}{1\ rev})^2\cdot0.02\ m}{9.8\ \dfrac{m}{s^2}}=0.72"